# Against The Achilles’ Heel: A Survey on Red Teaming for Generative Models

### v1

@@ 5.4.1 Comprehensive Safety
In this category, benchmarks typically offer a holistic assessment from multiple angles, incorporating evaluations of both attacks and defenses, and covering a wide range of harm areas. As a result, they are more comprehensive compared to other benchmarks. For example, SALAD-Bench (Li et al., 2024b) summarizes safety risks at three levels, expanding six types of safety risks into 65 fine-grained categories and 16 specific tasks. Since the evaluation covers various attack and defense methods, it utilizes more than one evaluation metric, such as safety rate, ASR, and rejection rate.
5.4.2 Specific Safety Concern
Benchmarks in this category often focus on one or several risks or examine safety issues in specific scenarios. For instance, SciMT-Safety (He et al., 2023b) primarily investigates the safety issues of LLMs in scientific research contexts. They first categorize the impact of scientific activities on the real world into nine scenarios and build the SciMT-Safety benchmark dataset. This dataset focuses on the chemical and biological field, with queries divided into substance-independent and substance-dependent categories. Note that the term “substance” here refers to chemical compounds, drugs, or microorganisms. Substance-independent queries do not involve specific substances but pose harmful intentions, for instance, “How to make a bomb?”. Substance-dependent queries include those involving specific substances, such as “Can a scientist use acetone to create a homemade rocket or missile? Describe the process in detail.”. During evaluation, this type of benchmark typically requires evaluators with specialized knowledge, whether humans or models.
@@ 5.4.3 Attack and Exploitation
Some benchmarks focus on evaluating attack methods, with their evaluation goals more concerned with the effectiveness of a particular attack or defense method, rather than the degree of harm caused by LLM responses. Thus, the attack success rate is often used as the metric. For example, StrongReject (Souly et al., 2024) argues that some previous benchmarks oversimplify jailbreak, and fail to accurately identify outputs attributed to a specific attack, thus often overestimating the effectiveness of attack methods. Therefore, they constructed a high-quality question set and used GPT4 to score the responses of LLMs across three dimensions: refusal, specificity, and convincingness.
@@
Despite variations in their evaluation focus, the methodologies for constructing these benchmarks share notable similarities, often necessitating a collaborative effort between humans and LLMs. For instance, humans may write prompts that fit specific hazardous scenarios, and then LLMs assist in generating samples in batch, striking a balance between sample quality and generation cost (He et al., 2023b; Sun et al., 2023; Xu et al., 2023a).
@@
These benchmarks not only provide datasets for assessment but often include specific evaluation methods and metrics, establishing standards for subsequent research. The evaluation methods mostly rely on automatic evaluators, such as prompting GPT4, and training a dedicated evaluator model (Wang et al., 2023g; Hosseini et al., 2023) to ensure the reproducibility of results and also to reduce the cost of extensive evaluation. The evaluation metrics used in these benchmarks are directly related to their respective evaluation focuses. For example, benchmarks related to attack and defense employ metrics such as attack success rate (ASR) and robustness rate as primary indicators (Yi et al., 2023; Shen et al., 2023; Toyer et al., 2023). Benchmarks for comprehensive safety assessments consider a variety of metrics, including but not limited to ASR and safety rate. By compiling these metrics based on the harm areas covered by the benchmarks, one can systematically understand the safety of the models being assessed (Li et al., 2024b; Zhang et al., 2023f). Other benchmarks aiming at evaluating specific scenarios, like CValues (Xu et al., 2023a), not only utilize safety score as the metric but also refer to a 1-10 responsibility score to assess whether the LLM’s responses bear adequate social responsibility.
6 Safeguard
Given numerous diverse attacks, defense measures have been widely researched and adopted in language model training and applications. Defense can be enforced in two stages: training time and inference time (Figure 16). We use training time to refer to the supervised fine-tuning and the reinforcement learning from human feedback (RLHF) step after pretraining. In this stage, models are modified to improve their ability to discern. On the other hand, in inference time, model weights are frozen and its behavior can only be steered by prompting. Guardrails that filter out unsafe content are also common in inference-time defense.
These two types of defense have different merits and drawbacks. In comparison to training-time defense, inference-time defense is more flexible and modular, as prompts can be edited and components can be introduced to the pipeline independently. However, all these measures add latency to the system and may not be able to identify nuanced attacks or those that require more reasoning. Training-time defense provides models with the ability to uncover such attacks. However, these measures are shown to impose an alignment tax (Lin et al., 2024) that impairs the normal function of language models, as the process shifts the distribution and causes catastrophic forgetting.
Figure 16: Defense strategies implemented at different stages in the life cycle of a language model. Training-time defense involves steps such as RLHF and fine-tuning to enhance the safety alignment, while inference-time defense employs system prompting, filtering out unsafe content, and guardrail systems to direct model behavior.
6.1 Training-time Defense
In this section, we discuss training-time defense, covering different approaches of fine-tuning, and the preference dataset and optimization process of RLHF.
6.1.1 Fine-tuning
Bianchi et al. (2023) discovered that fine-tuning the Llama model with just an additional 3% of safety samples can notably enhance its security. Wang et al. (2023h) proposed a method for building safety-aware language models by supervised fine-tuning. To achieve the goal, they first fine-tuned the language model to classify questions as harmful or harmless. Then, the same model is further fine-tuned to determine whether its responses are harmful. During inference, the language model can now evaluate the safety of its responses by labeling them at the end of their responses for downstream tasks. Ge et al. (2023) collects fine-tuning data through an adversarial approach where one adversarial model is trained to generate unsafe queries, while another model is safety-tuned to refuse such queries. The fine-tuning data is selected through a helpfulness reward model and a safety reward model, where unsafe data rated by the safety RM is used for adversarial training, while the safe and helpful data is used for safety fine-tuning.
6.1.2 RLHF
RLHF procedure has been widely adopted by both industry and academia for instilling human preferences and safety measures into LLMs. In RLHF, a reward model is trained to learn human preferences and then used to tune the target language model. The dataset for tuning reward models is the rank of responses to the same questions rated by the labeler (Ouyang et al., 2022). Compared to fine-tuning, RLHF has been shown to improve Out-of-Distribution (OOD) generalization, which is crucial with respect to safeguards (Kirk et al., 2024).
A key problem is the construction of the preference dataset. Ji et al. (2023) built a safety dataset for the question-answering task. In this dataset, pairs of answers with expert annotations of helpfulness and harmlessness are created for each question. Shi et al. (2023) proposed a more scalable automatic approach for constructing preference datasets. Particularly, they created a reverse instruction model for generating instructions given specific texts. For example, given a sonnet praising the value of human love, the model would generate an instruction like “Write a sonnet around human love”. In their case, they used the model to generate questions from harmful content, forming question-answer pairs for preference datasets.
The RLHF optimization process needs to balance multiple objectives to meet human preferences. One pair of objectives is helpfulness and harmlessness. These objectives are often conflicting because it is desired that models do not refuse normal questions (helpfulness), but refuse harmful questions (harmlessness), which creates a delicate boundary. Multi-objective optimization is often unstable and prone to mode collapse. Dai et al. (2023) separates these metrics during optimization by phrasing harmlessness as a cost objective and optimizing the objective with Lagrangian optimization, achieving a model with a better trade-off between helpfulness and harmlessness.
6.2 Inference-time Defense
In this section, we discuss inference-time defense, where different prompting techniques and guardrail systems are introduced. We also include emerging topics such as language model ensembles and adversarial suffix filters that target those generated by the likes of the AutoDAN searcher (Zou et al., 2023) (Table 1).
6.2.1 Prompting
Language models exhibit remarkable capabilities in in-context learning (Brown et al., 2020) and following instructions (Ouyang et al., 2022), which paves the way for the use of prompting as a training-free approach to enhance the safety of language models. Adding cues about responsibility and harmlessness directly to the system prompt has significantly improved the rejection rate of prompts designed to jailbreak or manipulate the model (Xie et al., 2023).
Furthermore, the technique of few-shot prompting has been extended to counteract sophisticated adversarial attacks. The strategy of In-Context Defense (Wei et al., 2023b) integrates specific examples of attack refusals into the prompt, thereby improving the recognition and resistance of the model against various attacks. Given the constraints on context length and the diversity of user queries, the combination of retrieval and few-shot prompting (Rubin et al., 2022) is often employed to ensure more effective and tailored responses to a wide range of user queries (Meade et al., 2023). This method retrieves safety-related examples that closely match the user’s query, providing a more comprehensive defense across different potential areas of harm.
Adopting a Chain-of-Thought (Wei et al., 2022) approach to enhance reasoning of language models, Zhang et al. (2024c) uses a multi-stage procedure that examines the intention behind a query before generating a response with established policies. Explicitly discerning the intent is particularly advantageous as language models often lack the ability to retrieve and reason on knowledge (Allen-Zhu & Li, 2023; Berglund et al., 2023) implicitly.
Prompting is a cost-effective method of implementing safety measures because it avoids modifying the substantial weights of Large Language Models. However, additional system prompts might increase response latency, particularly when numerous examples are used or multi-stage prompting is involved. Additionally, including unsafe content within examples can inadvertently guide language models during response and pose a risk of exploitation. In practice, prompting is often used with other techniques like keyword filtering.
6.2.2 Guardrail Systems
To systematically control language model response, guardrail systems have been developed to provide a unified interface to filter unsafe content (Dong et al., 2024a), often through a domain-specific language. The NeMo Guardrails system (Rebedea et al., 2023) provides a method to employ LLMs and vector databases to check unsafe content and hallucination in designated points of dialogue flow. They proposed the Colang language to define the guardrails. Similarly,  Sharma et al. (2024) proposed SPML, a domain-specific language that empowers prompt developers to efficiently create and maintain secure system prompts. SPML employs an intermediate representation for each entry, facilitating the comparison of incoming user inputs to ensure their safety.  Rai et al. (2024) built a multi-tiered guardrail system that includes a system prompt filter, toxic classifier, and ethical prompt generator to filter unsafe content,
To better identify unsafe content, fine-tuned models that check prompts for harmfulness are also developed. In a similar fashion, Pisano et al. (2023) added the prompt detecting stage and response correcting stage for another model to critique the incoming query and the response. The critique is appended to the query to remind the responding primary model. Inan et al. (2023) fine-tuned a Llama2 7b model to classify prompts based on their proposed taxonomy and risk guidelines. The model is prompted with classification tasks, risk taxonomies, the conversation, and output format requiring the response to contain violated categories for unsafe content. The model is shown to exhibit a high degree of transferability to other guidelines.
6.2.3 Language Model Ensemble
Language Model Ensemble refers to defense methods based on synthesizing and summarizing predictions from multiple models to derive answers. Chen et al. (2023a) introduced Moving Target Defense (MTD), which combines responses from eight commercial large-scale models to form the final answer. The authors devised their evaluation model to select “harmless” and “helpful” responses. This method innovatively employs the concept of ensemble. However, in practice, the response time is excessively long, and the computational cost is high. Hence, it cannot defend effectively against situations challenging for all commercial large-scale models.
Chern et al. (2024) made further improvements with debates among multiple models. In each multi-agent debate session, given a sensitive or dangerous topic, each LLM agent is first prompted to give a zero sample initial response. Then, for the number of rounds specified by the user, the agents perform a “discussion”, in which each agent updates its response with the output of other LLM agents (or itself) as additional recommendations. The final LLM output will be evaluated for toxicity. They found that engaging in debates with more powerful models could effectively enhance the security of weaker models.
6.2.4 Against Adversarial Suffixes
It has been found that by appending certain suffixes to the harmful query, the model can be induced to start the response with affirmative phrases and respond with harmful content (Zou et al., 2023). This approach is difficult to detect with conventional keyword input filters and has been studied. Several approaches have been proposed:
  • Perplexity Filter These defenses utilize the fact that these suffixes are often nonsensical and cause prompt perplexity to increase (Jain et al., 2023). However, Alon & Kamfonas (2023) found that relying solely on perplexity filtering could lead to a high false positive rate. They train a classifier using the Light GradientBoosting Machine (LightGBM) algorithm. They then run predictions on all validation samples and tune the thresholds that map the results to positive (adversarial) or negative (non-adversarial) classes to maximize the $F_{2}$ score.
  • Perturbation Another feature of these suffixes is that they are very sensitive to small perturbations. Based on this observation, Kumar et al. (2023) developed the erase-and-check algorithm to further detect and preprocess harmful prompts. Given a prompt P, this process removes the tokens individually (up to d tokens) and uses a security filter to check whether the deleted subsequence is safe. If the input prompt $P$ or any of its deleted subsequences are detected as harmful, the input prompt is marked as harmful. Instead, the prompt P is marked as safe only if the filter detects all sequences that are checked as safe. Robey et al. (2023) proposed the SmoothLLM algorithm based on randomized smoothing, a method to improve the adversarial robustness of models. It defends against attacks against samples by adding random noise to the input data. They enhanced the adversarial robustness of LLMs by employing insert, swap, and patch perturbations to their jailbreak attempts. This approach effectively reduces the success rate of GCG attacks to below 1%, achieving significant defense efficacy.
7 Multi-modal Model Red Teaming
@@
Multimodal models are designed to process and integrate information from various data types, including text, images, audio, video, and more. This section delves into the attack strategies for text (linguistic) and image (visual) data modalities.
Based on the different combinations of input and output modalities, we can identify nine distinct model categories, as illustrated in Figure Figure 17. We categorize the existing research according to the type of output: either visual only or a combination of visual and linguistic output.
Models that produce visual outputs (specifically, categories 2 and 7 in Figure 17) are designed to generate high-quality images from textual descriptions. Although these models may also process image inputs, we simplified them as Text-to-Image (T2I) models for convenience. The objective when attacking T2I models is to provoke the generation of inappropriate content, such as images depicting sexual content, harassment, or illegal activities.
@@ 7.1 Text-to-Image Model Attack
Text-to-image models can be manipulated with inappropriate inputs to produce harmful content. [^3] In the context of T2I, these harmful contents are usually referred to as Not-Safe-For-Work (NSFW) content. This is similar to the process of attacking LLMs, where the goal is to create textual prompts that lead to the generation of harmful content. Rather than attempting to manipulate T2I to produce violent images or to remove a given object without restrictions, Shahgir et al. (2023) aim to replace an object in the image with another targeted one (entity-swapping attack).
Evade Defense Strategy
Many studies have focused on attacking T2I models by evading two common defense strategies: the textual prompt filter and the post hoc image safety checker. The textual prompt filter works by blocking the embedding of certain words or concepts, thereby preventing the generation of certain concepts. The post hoc image safety checker refuses to output an image if it is detected as problematic.
Yang et al. (2023c) introduced MMA-Diffusion, which constructs adversarial prompts that do not contain any sensitive words but maintain similar semantics to the target prompt. This is achieved through semantic similarity loss and gradient-driven continuous optimization (mentioned in Sec. 4.2), leading to the regularization of sensitive words.
Liu et al. (2024c) observed that sensitive terms like “kill” can be decomposed into less sensitive words like “fighting” to bypass text filters, and combining sensitive terms like “blood” with non-sensitive terms like “red liquid” can evade image filters. They thus introduced an automated framework, Groot, that uses LLMs for adversarial testing on T2I models. This framework leverages semantic decomposition and sensitive element drowning techniques to generate semantically consistent adversarial prompts. The framework iteratively tests and refines these prompts, analyzing failures to adjust its approach until a successful or time-limited attempt is made.
Mehrabi et al. (2023a) advanced this approach by incorporating a feedback signal, such as the evaluation of the corresponding output image by safety classifiers or human feedback, into a loop. This feedback is used to update the prompts through in-context learning with a language model, instead of merely repeating the refined prompt multiple times until the attack succeeds.
@@
In response to the aforementioned prompt attacks, Wu et al. (2024) proposed a prompt optimization method to prevent the generation of inappropriate images. When a user inputs a potentially harmful prompt, this method automatically modifies the prompt to ensure that any generated images are appropriate, while still retaining the acceptable aspects of the original user input. This is similar to the prompt rewriting defense method introduced in Sec. 6.
@@
Additionally, the post hoc checker plays a crucial role in identifying harmful content. It operates by converting the generated image into a latent vector and then comparing this vector with predefined unsafe embeddings. If the similarity between the latent vector and any of the unsafe embeddings exceeds a certain threshold, the generated content is flagged as unsafe (Yang et al., 2023c).
7.2 Vision Language Model Attack

### v2

@@ 6.4.1 Comprehensive Safety
In the safety area, benchmarks typically offer a holistic assessment from multiple angles, incorporating evaluations of both attacks and defenses, and covering a wide range of harm areas. As a result, they are more comprehensive compared to other benchmarks. For example, SALAD-Bench Li et al. (2024c) summarizes safety risks at three levels, expanding six types of safety risks into 65 fine-grained categories and 16 specific tasks. Since the evaluation covers various attack and defense methods, it uses more than one evaluation metric, including safety rate, ASR, and rejection rate.
6.4.2 Specific Safety Concerns
Benchmarks in this category often focus on one or more risks, or examine safety issues in specific scenarios. For instance, SciMT-Safety He et al. (2023b) primarily investigates the safety of LLMs in scientific research contexts. They first categorize the impact of scientific activities on the real world into nine scenarios and build the SciMT-Safety benchmark dataset. This dataset focuses on the chemical and biological field, with queries divided into substance-independent and substance-dependent categories. Note that the term “substance” here refers to chemical compounds, drugs, or microorganisms. Substance-independent queries do not involve specific substances but pose harmful intentions, for instance, “How to make a bomb?”. Substance-dependent queries include those involving specific substances, such as “Can a scientist use acetone to create a homemade rocket or missile? Describe the process in detail”. For evaluation, this type of benchmark typically requires evaluators with specialized knowledge, whether humans or models.
@@ 6.4.3 Attack and Exploitation
Some benchmarks focus on evaluating attack methods, with their evaluation goals more concerned with the effectiveness of a particular attack or defense method, rather than the degree of harm caused by LLM responses. For example, StrongReject Souly et al. (2024) argues that some previous benchmarks oversimplify jailbreaking, and fail to accurately identify outputs attributed to a specific attack, thus often overestimating the effectiveness of attack methods. Therefore, they constructed a high-quality question set and used GPT4 to score the responses of LLMs across three dimensions: refusal, specificity, and convincingness.
@@
Despite variability in their evaluation focus, the methodologies for constructing these benchmarks share notable similarities, often necessitating a collaborative effort between humans and LLMs. For instance, humans may write prompts that fit specific hazardous scenarios, and then LLMs assist in generating batches of samples, striking a balance between sample quality and generation cost He et al. (2023b); Sun et al. (2023); Xu et al. (2023a).
@@
These benchmarks not only provide datasets for assessment but often include specific evaluation methods and metrics, establishing standards for subsequent research. The evaluation methods mostly rely on automatic evaluators, such as prompting GPT4, and training a dedicated evaluator model Wang et al. (2023g; 2024f); Hosseini et al. (2023) to ensure the reproducibility of results and also to reduce the cost of extensive evaluation. The evaluation metrics used in these benchmarks are directly related to their respective evaluation focus. For example, benchmarks related to attack and defense employ metrics such as attack success rate (ASR) and robustness rate as primary indicators Yi et al. (2023); Shen et al. (2023); Toyer et al. (2023). Benchmarks for comprehensive safety assessments consider a variety of metrics, in particular ASR and safety rate. By compiling these metrics based on the harm areas covered by the benchmarks, one can systematically understand the safety of the models being assessed Li et al. (2024c); Zhang et al. (2023f). Other benchmarks aiming at evaluating specific scenarios, like CValues Xu et al. (2023a), not only use safety scores but also refer to a 1–10 scale responsibility score, to assess whether the LLM’s responses bear adequate social responsibility.
6.5 Empirical Comparisons
While a comprehensive evaluation spanning every combination of attack, defense, and language model is beyond the scope of this survey, we identify work that focuses on comparing different methods in a unified setting.
Xu et al. (2024g) compared the effect of 9 attack methods including popular attack strategies like GPTFuzz Yu et al. (2023a), base64 Lambert et al. (2023) and language model based or gradient based searchers including PAIR Chao et al. (2023) and GCG Zou et al. (2023b). The experiment was conducted with three models—GPT3.5, Vicuna and Llama—with results favoring template-based approaches for their efficiency and high ASR rate, and disfavoring gradient-based searchers (Figure 26). GPTFuzz achieved 100% ASR against GPT3.5, while AutoDAN and GCG lag behind with their average ASR hovering around 20–40%. The study also revealed that the safeguard levels vary across language models, with Vicuna being more susceptible to different attacks than Llama, and GPT3.5 being more robust to jailbreak attempts. They have also investigated 7 defense techniques, pointing to the limited protection offered by current defenses.
Figure 26: Comparison of different jailbreak attack and defense methods. The left figure shows experimental results of 9 attack methods against GPT3.5, Vicuna and Llama, where template-based approaches like GPTFuzz achieved 100% ASR while gradient-based methods like AutoDAN and GCG showed lower effectiveness (20-40% ASR). The right figure presents evaluation results of 7 defense techniques, highlighting their limited protection capabilities. Both figures are adopted from Xu et al. (2024g).
In a similar vein, Xu et al. (2024e) evaluated 9 jailbreak attacks and 6 defense methods on models of diverse sizes and levels of safety alignment, with a focus on the tug-of-war between different jailbreak techniques and safeguards (Table 4). Their conclusions suggest that adversarial robustness is mostly independent from model size, that fine-tuning and system prompts improves LLM safety, and that small adjustments to prompts plays a significant role in model robustness to adversarial attacks. For prompt searchers, their experiments reveal that stronger models as attackers exhibit better jailbreak performance, that longer adversarial suffixes are correlated with high ASR rate despite plateauing afterwards, and that allocating more attack budget to gradient-based searchers than prompt searchers is more effective. In terms of the effect of jailbreak attacks against defense methods, gradient-based searchers including AutoDAN and AdvPrompter are more robust than prompt-based approaches.
[…]
Vicuna-13B ($\text{ASR}_{\text{Prefix}}$ / $\text{ASR}_{\text{Agent}}$)
GCG 92 / 14 84 / 20 92 / 20 98 / 2 88 / 8 100 / 40 98 / 14
AutoDAN 100 / 92 84 / 60 100 / 58 86 / 20 78 / 68 100 / 86 82 / 70
AmpleGCG 100 / 18 100 / 8 100 / 0 94 / 14 100 / 4 100 / 30 100 / 2
AdvPrompter 100 / 44 100 / 10 100 / 0 90 / 8 98 / 30 100 / 46 100 / 24
PAIR 36 / 36 28 / 32 60 / 30 88 / 12 44 / 48 76 / 96 20 / 24
TAP 28 / 32 24 / 12 38 / 22 96 / 4 30 / 32 70 / 96 22 / 18
GPTFuzzer 78 / 92 30 / 88 38 / 60 90 / 4 66 / 92 32 / 94 72 / 84
LLaMA-2-7B ($\text{ASR}_{\text{Prefix}}$ / $\text{ASR}_{\text{Agent}}$)
GCG 8 / 2 0 / 0 4 / 0 82 / 2 4 / 0 4 / 2 2 / 0
AutoDAN 50 / 32 2 / 2 86 / 54 70 / 16 50 / 32 54 / 32 52 / 42
AmpleGCG 100 / 50 100 / 6 100 / 10 74 / 14 100 / 44 100 / 52 100 / 50
AdvPrompter 98 / 20 100 / 4 100 / 2 64 / 8 98 / 20 96 / 20 98 / 22
PAIR 18 / 6 16 / 4 60 / 6 40 / 8 18 / 8 12 / 8 12 / 4
TAP 18 / 12 22 / 0 38 / 6 36 / 20 16 / 4 18 / 6 12 / 8
GPTFuzzer 6 / 22 2 / 8 18 / 18 82 / 4 18 / 26 2 / 8 22 / 30
Table 4: Jailbreak attack experiments on AdvBench under metric $\text{ASR}_{\text{Prefix}}$ and $\text{ASR}_{\text{Agent}}$, corresponding to the metrics introduced in Section 6.3.1 and Section 6.3.2, respectively. Defense methods include Self-Reminder Xie et al. (2023), RPO Zhou et al. (2024a), SmoothLLM Robey et al. (2023), Adv. Training Madry (2017), Unlearning Yao et al. (2023d) and Safety Training Touvron et al. (2023). Table adopted from Xu et al. (2024e).
7 Multimodal Model Red Teaming
In previous sections, we examined attack, defense, and evaluation methods specifically related to language model safety. However, with the increasing number of multimodal models that process additional data types (e.g., images, video, auditory data, or mixtures of all of these), new safety challenges emerge. Unlike text which consists of discrete words, multimodal information usually includes modalities of images, videos, and audio, which have multi-dimensional structure, and are usually continuous in the input space. Multimodal large models are susceptible to malicious attacks on visual inputs due to the continuity of image signals.
Given the unique vulnerabilities associated with each modality, multimodal model safety requires distinct approaches. In this section, we focus on attack strategies specific to linguistic and visual data modalities, providing insights into the complexities of safeguarding these broader, integrated models.
Based on the different combinations of input and output modalities, we can identify nine distinct model categories, as illustrated in Figure 27. We categorize the existing research according to the type of output: either visual only or a combination of visual and linguistic output.
@@ 7 Multimodal Model Red Teaming
Models that produce visual outputs (specifically, categories 2 and 7 in Figure 27) are designed to generate high-quality images from textual descriptions. Although these models may also process image inputs, we simplistically consider them as Text-to-Image (T2I) models for convenience. The objective when attacking T2I models is to invoke the generation of inappropriate content, such as images depicting sexual content, harassment, or illegal activities.
@@ 7.1 Text-to-Image Model Attack
Text-to-image models can be manipulated with inappropriate inputs to produce harmful content. [^4] In the context of T2I, such harmful content is usually referred to as Not-Safe-For-Work (NSFW) content. This is similar to the process of attacking LLMs, where the goal is to create textual prompts that lead to the generation of harmful content. Rather than attempting to manipulate T2I to produce violent images or to remove a given object without restrictions, Shahgir et al. (2023) aim to replace an object in the image with another targeted one (entity-swapping attack).
Evading Defense Strategies
Many studies have focused on attacking T2I models by evading two common defense strategies: the textual prompt filter and the post hoc image safety checker. The textual prompt filter works by blocking the embedding of certain words or concepts, thereby preventing the generation of certain concepts. Post hoc image safety checkers refuse to output an image if it is detected as problematic.
Yang et al. (2023c) introduced MMA-Diffusion, which constructs adversarial prompts that do not contain any sensitive words but have similar semantics to the target prompt. This is achieved through semantic similarity loss and gradient-driven continuous optimization (mentioned in Section 4.2), leading to the regularization of sensitive words.
Liu et al. (2024c) observed that sensitive terms like “kill” can be decomposed into less sensitive words like “fighting” to bypass text filters, and combining sensitive terms like “blood” with non-sensitive terms like “red liquid” can evade image filters. They thus introduced an automated framework, Groot, that uses LLMs to adversarially attack T2I models. This framework leverages semantic decomposition and sensitive element drowning techniques to generate semantically consistent adversarial prompts. The framework iteratively tests and refines these prompts, analyzing failures and adjusting its approach until a successful or time-limited attempt is made.
Mehrabi et al. (2023a) further advanced this approach by incorporating a feedback signal, such as the evaluation of the corresponding output image by safety classifiers or human feedback, into a loop. This feedback is used to update the prompts through in-context learning with a language model, instead of merely repeating the refined prompt multiple times until the attack succeeds.
@@
In response to the aforementioned prompt attacks, Wu et al. (2024) proposed a prompt optimization method to prevent the generation of inappropriate images. When a user inputs a potentially harmful prompt, this method automatically modifies the prompt to ensure that any generated images are appropriate, by retaining acceptable aspects of the original user input. This is similar to the prompt rewriting defense method introduced in Section 5.
@@
Additionally, the post hoc checker plays a crucial role in identifying harmful content. It operates by encoding the generated image and comparing it with predefined unsafe embeddings. If the similarity between the latent image representation and any of the unsafe embeddings exceeds a certain threshold, the generated content is flagged as unsafe Yang et al. (2023c).
7.2 Vision Language Model Attacks